OpenStudy (anonymous):

Solve the interval [0, 2pi) 2sinTheta^2-cosTheta-1=0

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Can you guide me in the direction of solving this equation? \[2\sin \theta-\cos \theta-1=0\]

OpenStudy (ipwnbunnies):

Is it:\[2\sin^{2}\Theta - \cos \Theta - 1 = 0\]

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

Consider the equation: cos (2 theta ) = sin (theta)

OpenStudy (anonymous):

yes it is \[2\sin \theta ^{2}-\cos \theta-1=0\]

OpenStudy (anonymous):

Using identity cos (2 theta ) = 1 - 2 ( sin (theta) )^2, the above equation becomes: 1 - 2 ( sin (theta) )^2 = sin(theta) which can be rewritten as 2 ( sin (theta) )^2 + sin(theta) -1 = 0 Now this is is a quadratic equation in one variable .Either solve it using the formula for the solutions to a quadratic equation OR use the following method: Above equation can be rewritten as ( sin(theta) + 1 ) ( 2sin(theta) -1 ) = 0 Thus we have sin(theta)= -1 OR sin(theta)= 0.5 Hence, theta=pi/6, 5pi/6 OR 3pi/2 That is all.

OpenStudy (anonymous):

you get it.

OpenStudy (anonymous):

Thank you, I was wondering If I could solve it using my identities with or without switching to the RHS. Thank you for clearing it up for me! :)

OpenStudy (anonymous):

no prob